OpenStudy (anonymous):

Use the information given about the angle θ, 0≤θ≤2π, to find the exact value of the indicated trigonometric function. cscθ=-5/2, tanθ>0, find cos θ/2.

OpenStudy (loser66):

are you Vietnamese?

OpenStudy (zpupster):

csc=hyp/opp 1/opp/hyp = 1/sin= 1/-5/2 or 2/-5 us pythagorean to adjacent side hyp^2 - Opp^2= adj^2 or sqrt(nyp^2-opp^2) = adj

OpenStudy (anonymous):

@Loser66: yes :) @zpupster: thanks

OpenStudy (loser66):

Can you read Vietnamese? Chi la nguoi Viet day. Neu em can hoi gi, neu biet chi se uu tien cho. hehehe

OpenStudy (zpupster):

You’re Welcome = Không Có Chi No Problem = Không Sao